# The Diophantine problem in isotropic reductive groups

### v1

@@ 2 Isotropic reductive groups
Exceptional Tits indices are given in the following table. Sizes and types of preimages are given by increasing the root length. The indices of labels in $J$ denote the lengths of the corresponding roots in $\Phi$, namely, long, short, and ultrashort ones.
Tits index $\Phi$ $J$ $u^{-1}(0)$ $|u^{-1}(\alpha)|$ $u^{-1}(\mathbb{Z}\alpha\cap(\Phi\cup\{0\}))$
${{}^{3}\!{\mathsf{D}_{4,0}^{28}}}$, ${{}^{6}\!{\mathsf{D}_{4,0}^{28}}}$ 00 $\varnothing$ $\mathsf{D}_{4}$
${{}^{1}\!{\mathsf{E}_{6,0}^{78}}}$, ${{}^{2}\!{\mathsf{E}_{6,0}^{78}}}$ 00 $\varnothing$ $\mathsf{E}_{6}$
$\mathsf{E}_{7,0}^{133}$ 00 $\varnothing$ $\mathsf{E}_{7}$
$\mathsf{E}_{8,0}^{248}$ 00 $\varnothing$ $\mathsf{E}_{8}$
$\mathsf{F}_{4,0}^{52}$ 00 $\varnothing$ $\mathsf{F}_{4}$
$\mathsf{G}_{2,0}^{14}$ 00 $\varnothing$ $\mathsf{G}_{2}$
$\mathsf{E}_{7,1}^{78}$ $\mathsf{A}_{1}$$\{7\}$ $\mathsf{E}_{6}$ $27$ $\mathsf{E}_{7}$
${{}^{3}\!{\mathsf{D}_{4,1}^{9}}}$, ${{}^{6}\!{\mathsf{D}_{4,1}^{9}}}$ $\mathsf{BC}_{1}$$\{2\}$ $3\mathsf{A}_{1}$ $8$, $1$ $\mathsf{D}_{4}$, $4\mathsf{A}_{1}$
${{}^{2}\!{\mathsf{E}_{6,1}^{35}}}$ $\mathsf{BC}_{1}$$\{2\}$ $\mathsf{A}_{5}$ $20$, $1$ $\mathsf{E}_{6}$, $\mathsf{A}_{1}+\mathsf{A}_{5}$
$\mathsf{E}_{7,1}^{66}$ $\mathsf{BC}_{1}$$\{1\}$ $\mathsf{D}_{6}$ $32$, $1$ $\mathsf{E}_{7}$, $\mathsf{A}_{1}+\mathsf{D}_{6}$
$\mathsf{E}_{8,1}^{133}$ $\mathsf{BC}_{1}$$\{8\}$ $\mathsf{E}_{7}$ $56$, $1$ $\mathsf{E}_{8}$, $\mathsf{A}_{1}+\mathsf{E}_{7}$
$\mathsf{F}_{4,1}^{21}$ $\mathsf{BC}_{1}$$\{4\}$ $\mathsf{B}_{3}$ $8$, $7$ $\mathsf{F}_{4}$, $\mathsf{B}_{4}$
${{}^{2}\!{\mathsf{E}_{6,1}^{29}}}$ $\mathsf{BC}_{1}$$\{1,6\}$ $\mathsf{D}_{4}$ $16$, $8$ $\mathsf{E}_{6}$, $\mathsf{D}_{5}$
$\mathsf{E}_{7,1}^{48}$ $\mathsf{BC}_{1}$$\{6\}$ $\mathsf{A}_{1}+\mathsf{D}_{5}$ $32$, $10$ $\mathsf{E}_{7}$, $\mathsf{D}_{6}$
$\mathsf{E}_{8,1}^{91}$ $\mathsf{BC}_{1}$$\{1\}$ $\mathsf{D}_{7}$ $64$, $14$ $\mathsf{E}_{8}$, $\mathsf{D}_{8}$
${{}^{1}\!{\mathsf{E}_{6,2}^{28}}}$ $\mathsf{A}_{2}$$\{1,6\}$ $\mathsf{D}_{4}$ $8$ $\mathsf{D}_{5}$
$\mathsf{G}_{2,2}^{0}$ $\mathsf{G}_{2}$$\{1_{\mathrm{s}},2_{\mathrm{l}}\}$varnothing $1$, $1$ $\mathsf{A}_{1}$, $\mathsf{A}_{1}$
${{}^{3}\!{\mathsf{D}_{4,2}^{2}}}$, ${{}^{6}\!{\mathsf{D}_{4,2}^{2}}}$ $\mathsf{G}_{2}$$\{1_{\mathrm{s}},2_{\mathrm{l}},3_{\mathrm{s}},4_{\mathrm{s}}\}$varnothing $3$, $1$ $3\mathsf{A}_{1}$, $\mathsf{A}_{1}$
${{}^{1}\!{\mathsf{E}_{6,2}^{16}}}$, ${{}^{2}\!{\mathsf{E}_{6,2}^{16^{\prime\prime}}}}$ $\mathsf{G}_{2}$$\{2_{\mathrm{l}},4_{\mathrm{s}}\}$ $2\mathsf{A}_{2}$ $9$, $1$ $\mathsf{A}_{5}$, $\mathsf{A}_{1}+2\mathsf{A}_{2}$
$\mathsf{E}_{8,2}^{72}$ $\mathsf{G}_{2}$$\{7_{\mathrm{s}},8_{\mathrm{l}}\}$ $\mathsf{E}_{6}$ $27$, $1$ $\mathsf{E}_{7}$, $\mathsf{A}_{1}+\mathsf{E}_{6}$
${{}^{2}\!{\mathsf{E}_{6,2}^{16^{\prime}}}}$ $\mathsf{BC}_{2}$$\{1_{\mathrm{us}},2_{\mathrm{s}},6_{\mathrm{us}}\}$ $\mathsf{A}_{3}$ $8$, $6$, $1$ $\mathsf{A}_{5}$, $\mathsf{D}_{4}$, $2\mathsf{A}_{1}+\mathsf{A}_{3}$
$\mathsf{E}_{7,2}^{31}$ $\mathsf{BC}_{2}$$\{1_{\mathrm{s}},6_{\mathrm{us}}\}$ $\mathsf{A}_{1}+\mathsf{D}_{4}$ $16$, $8$, $1$ $\mathsf{D}_{6}$, $\mathsf{A}_{1}+\mathsf{D}_{5}$, $2\mathsf{A}_{1}+\mathsf{D}_{4}$
$\mathsf{E}_{8,2}^{66}$ $\mathsf{BC}_{2}$$\{1_{\mathrm{us}},8_{\mathrm{s}}\}$ $\mathsf{D}_{6}$ $32$, $12$, $1$ $\mathsf{E}_{7}$, $\mathsf{D}_{7}$, $\mathsf{A}_{1}+\mathsf{D}_{6}$
$\mathsf{E}_{7,3}^{28}$ $\mathsf{C}_{3}$$\{1_{\mathrm{s}},6_{\mathrm{s}},7_{\mathrm{l}}\}$ $\mathsf{D}_{4}$ $8$, $1$ $\mathsf{D}_{5}$, $\mathsf{A}_{1}+\mathsf{D}_{4}$
$\mathsf{F}_{4,4}^{0}$ $\mathsf{F}_{4}$$\{1_{\mathrm{l}},2_{\mathrm{l}},3_{\mathrm{s}},4_{\mathrm{s}}\}$varnothing $1$, $1$ $\mathsf{A}_{1}$, $\mathsf{A}_{1}$
${{}^{2}\!{\mathsf{E}_{6,4}^{2}}}$ $\mathsf{F}_{4}$$\{1_{\mathrm{s}},2_{\mathrm{l}},3_{\mathrm{s}},4_{\mathrm{l}},5_{\mathrm{s}},6_{\mathrm{s}}\}$varnothing $2$, $1$ $2\mathsf{A}_{1}$, $\mathsf{A}_{1}$
$\mathsf{E}_{7,4}^{9}$ $\mathsf{F}_{4}$$\{1_{\mathrm{l}},3_{\mathrm{l}},4_{\mathrm{s}},6_{\mathrm{s}}\}$ $3\mathsf{A}_{1}$ $4$, $1$ $\mathsf{A}_{1}+\mathsf{A}_{3}$, $4\mathsf{A}_{1}$
$\mathsf{E}_{8,4}^{28}$ $\mathsf{F}_{4}$$\{1_{\mathrm{s}},6_{\mathrm{s}},7_{\mathrm{l}},8_{\mathrm{l}}\}$ $\mathsf{D}_{4}$ $8$, $1$ $\mathsf{D}_{5}$, $\mathsf{A}_{1}+\mathsf{D}_{4}$
${{}^{1}\!{\mathsf{E}_{6,6}^{0}}}$ $\mathsf{E}_{6}$$\{1,2,3,4,5,6\}$varnothing $1$ $\mathsf{A}_{1}$
$\mathsf{E}_{7,7}^{0}$ $\mathsf{E}_{7}$$\{1,2,3,4,5,6,7\}$varnothing $1$ $\mathsf{A}_{1}$
$\mathsf{E}_{8,8}^{0}$ $\mathsf{E}_{8}$$\{1,2,3,4,5,6,7,8\}$varnothing $1$ $\mathsf{A}_{1}$
@@ 3 Technical lemmas
Note that any parabolic set is saturated (and every saturated set is an intersection of parabolic ones), but $\mathsf{A}_{2}\subseteq\mathsf{G}_{2}$ is not saturated. Any closed subset $\Sigma$ admits a unique decomposition $\Sigma=\Sigma_{\mathrm{r}}\sqcup\Sigma_{\mathrm{u}}$, where $\Sigma_{\mathrm{r}}=\Sigma\cap(-\Sigma)$ is a closed root subsystem and $\Sigma_{\mathrm{u}}=\Sigma\setminus(-\Sigma)$ is a unipotent set. Conversely, if $\Sigma_{\mathrm{r}}$ is a closed root subsystem, $\Sigma_{\mathrm{u}}$ is a unipotent set disjoint with $\Sigma_{\mathrm{r}}$, and $(\Sigma_{\mathrm{r}}+\Sigma_{\mathrm{u}})\cap\Phi\subseteq\Sigma_{\mathrm{u}}$, then $\Sigma_{\mathrm{r}}\cup\Sigma_{\mathrm{u}}$ is a closed set. A closed set $\Sigma$ is parabolic if and only if there are $w\in\mathrm{W}(\Phi)$ and $J\subseteq\Delta$ such that $w\Sigma=(\Pi+\mathbb{Z}J)\cap\Phi$. Here $\Pi\subseteq\Phi$ is the set of positive roots and $\Delta\subseteq\Pi$ is the set of basic roots. The classes of closed sets, unipotent sets, closed root subsystems, and saturated sets are closed under intersection.
@@
If $\Sigma\subseteq\Phi$ is a closed root subsystem, then the smallest closed set containing $u^{-1}(\Sigma)$ consists of $u^{-1}(\Sigma)$ and sums of pairs of roots of $u^{-1}(\Sigma)$ (with opposite images in $\Sigma$), it is also a closed root subsystem.
For any closed subset $\Sigma\subseteq\Phi$ let $G_{\Sigma}\leq G$ and $G^{0}_{\Sigma}\leq G$ be the group subsheaves generated by $\bigcup_{\alpha\in\Sigma}U_{\alpha}$ and $L\cup\bigcup_{\alpha\in\Sigma}U_{\alpha}$, they are actually smooth closed group subschemes with connected geometric fibers over $K$ and $G^{0}_{\Sigma}$ is of type (RC) [3, §XXII.5]. More precisely, if $\Sigma$ is a root subsystem, then $G_{\Sigma}$ and $G^{0}_{\Sigma}$ are reductive group subschemes (with induced isotropic structure). If $\Sigma$ is unipotent, then $G_{\Sigma}$ is a unipotent subgroup, i.e. it is an étale twisted form of some $\mathbb{A}^{N}$ as a scheme. Finally, in general $G_{\Sigma}=G_{\Sigma_{\mathrm{r}}}\rtimes G_{\Sigma_{\mathrm{u}}}$, and $G^{0}_{\Sigma}=G^{0}_{\Sigma_{\mathrm{r}}}\rtimes G_{\Sigma_{\mathrm{u}}}$. If $\Sigma$ is parabolic, then $G^{0}_{\Sigma}$ is a parabolic group subscheme. For any parabolic $\Sigma\subseteq\Phi$ the multiplication morphism
@@
$G^{0}_{\Sigma}(K)=G_{\Sigma_{\mathrm{r}}\cap\Pi}(K)\,G_{\Sigma\cap(-\Pi)}(K)\,G_{\Sigma\cap\Pi}(K)\,L(K).$
@@
We may assume that $\operatorname{Spec}(K)$ is connected. In this case there is a maximal isotropic structure $(L^{\prime},G^{\prime}_{\alpha})_{\alpha\in\Phi^{\prime}}$ and a map $v\colon\Phi^{\prime}\to\Phi\cup\{0\}$ such that $\Phi\subseteq v(\Phi^{\prime})$, $G^{\prime}_{\alpha}=G_{v^{-1}(\alpha)}$, $L^{\prime}=G_{\Phi\cap v^{-1}(0)}$ [3, XXVI.7.4.2]. Gauss decomposition holds for $G$ with respect to a maximal isotropic structure by [3, corollary XXVI.5.2], i.e.
$G(K)=G^{\prime}_{\Pi^{\prime}}(K)\,G^{\prime}_{-\Pi^{\prime}}(K)\,G^{\prime}_{\Pi^{\prime}}(K)\,L^{\prime}(K).$
@@
$G^{0}_{\Sigma}(K)=G_{\Sigma\cap\Pi}(K)\,G_{\Sigma\cap(-\Pi)}(K)\,G_{\Sigma\cap\Pi}(K)\,L(K)$
@@
$G^{0}_{\Sigma}(K)$ $=G_{\Sigma_{\mathrm{r}}\cap\Pi}(K)\,G_{\Sigma_{\mathrm{r}}\cap(-\Pi)}(K)\,G_{\Sigma_{\mathrm{u}}}(K)\,G_{\Sigma_{\mathrm{r}}\cap\Pi}(K)\,L(K)$
$=G_{\Sigma_{\mathrm{r}}\cap\Pi}(K)\,G_{\Sigma\cap(-\Pi)}(K)\,G_{\Sigma\cap\Pi}(K)\,L(K).\qed$
For any closed subsets $\Sigma,\Sigma^{\prime}\subseteq\Phi$ the intersection $G^{0}_{\Sigma}\cap G^{0}_{\Sigma^{\prime}}$ is a smooth closed subscheme and its fiberwise connected component is $G^{0}_{\Sigma\cap\Sigma^{\prime}}$ by [3, proposition XXII.5.4.5].
@@
For the first claim it suffices to prove that $G^{0}_{\Sigma}\cap G^{0}_{\Sigma^{\prime}}=G^{0}_{\Sigma\cap\Sigma^{\prime}}$ if $\Sigma^{\prime}\supseteq\Pi$ is parabolic and $K$ is local. Take $g\in G^{0}_{\Sigma}(K)\cap G^{0}_{\Sigma^{\prime}}(K)$. By lemma 1 we may multiply $g$ from both sides by elements of $G^{0}_{\Sigma\cap\Sigma^{\prime}}$ to get a new element $g^{\prime}\in G_{\Sigma\cap(-\Sigma^{\prime}_{\mathrm{u}})}(K)\cap G^{0}_{\Sigma^{\prime}}(K)$. But such $g^{\prime}$ is necessarily trivial.
Now let us prove the second claim. Without loss of generality $\Sigma^{\prime}\subseteq\Pi$, so
$G^{0}_{\Sigma}\cap G_{\Sigma^{\prime}}=G^{0}_{\Sigma}\cap G^{0}_{\Pi}\cap G_{\Sigma^{\prime}}=G^{0}_{\Sigma\cap\Pi}\cap G_{\Sigma^{\prime}}=G_{\Sigma\cap\Pi\cap\Sigma^{\prime}}=G_{\Sigma\cap\Sigma^{\prime}}.\qed$
Recall [8, §4] that a $2$-step $K$-module $(M,M_{0})$ consists of
@@
  • a central subgroup $M_{0}\leq M$ such that $[M,M]^{\cdot}\leq M_{0}$;
@@
  • a right action $({-})\cdot({=})\colon M\times K\to M$ of the multiplicative monoid by group automorphisms such that $[m\cdot k,m^{\prime}\cdot k^{\prime}]^{\cdot}=kk^{\prime}[m,m^{\prime}]^{\cdot}$, $m\cdot(k+k^{\prime})=m\cdot k\dotplus kk^{\prime}\tau(m)\dotplus m\cdot k^{\prime}$ for some (uniquely determined) $\tau(m)\in M_{0}$, $m_{0}\cdot k=k^{2}m_{0}$ for $m_{0}\in M_{0}$.
@@
$\tau(\mathbin{\dot{-}}m)$ =˙$=\mathbin{\dot{-}}\tau(m),$ $\tau(m\cdot k)$ $=k^{2}\tau(m),$

### v2

@@ 2 Isotropic reductive groups
Exceptional isotropic irreducible Tits indices are given in the following table. Sizes and types of preimages are given by increasing the root length. The indices of labels in $J$ denote the lengths of the corresponding roots in $\Phi$, namely, long, short, and ultrashort ones. When working with these group schemes it is useful to use explicit root diagrams of all exceptional root systems [22], such diagrams split into the subdiagram $u^{-1}(0)$ and blocks $u^{-1}(\alpha)$ connected by edges and outer automorphisms of the Tits index.
Tits index $\widetilde{\Phi}$ $J$ $\Phi$ $u^{-1}(0)$ $|u^{-1}(\alpha)|$ $|\mathrm{Out}|$
$\mathsf{E}_{7,1}^{78}$ $\mathsf{E}_{7}$ $\{7\}$ $\mathsf{A}_{1}$ $\mathsf{E}_{6}$ $27$ 11
$^{3}\!{\mathsf{D}_{4,1}^{9}}$}, $^{6}\!{\mathsf{D}_{4,1}^{9}}$} $\mathsf{D}_{4}$ $\{2\}$ $\mathsf{BC}_{1}$ $3\mathsf{A}_{1}$ 88, 11 66
$^{2}\!{\mathsf{E}_{6,1}^{35}}$} $\mathsf{E}_{6}$ $\{2\}$ $\mathsf{BC}_{1}$ $\mathsf{A}_{5}$ 2020, 11 22
$\mathsf{E}_{7,1}^{66}$ $\mathsf{E}_{7}$ $\{1\}$ $\mathsf{BC}_{1}$ $\mathsf{D}_{6}$ $32$, 11 11
$\mathsf{E}_{8,1}^{133}$ $\mathsf{E}_{8}$ $\{8\}$ $\mathsf{BC}_{1}$ $\mathsf{E}_{7}$ $56$, 11 11
$\mathsf{F}_{4,1}^{21}$ $\mathsf{F}_{4}$ $\{4\}$ $\mathsf{BC}_{1}$ $\mathsf{B}_{3}$ 88, 77 11
$^{2}\!{\mathsf{E}_{6,1}^{29}}$} $\mathsf{E}_{6}$ $\{1,6\}$ $\mathsf{BC}_{1}$ $\mathsf{D}_{4}$ $16$, 88 22
$\mathsf{E}_{7,1}^{48}$ $\mathsf{E}_{7}$ $\{6\}$ $\mathsf{BC}_{1}$ $\mathsf{A}_{1}+\mathsf{D}_{5}$ $32$, $10$ 11
$\mathsf{E}_{8,1}^{91}$ $\mathsf{E}_{8}$ $\{1\}$ $\mathsf{BC}_{1}$ $\mathsf{D}_{7}$ $64$, $14$ 11
$^{1}\!{\mathsf{E}_{6,2}^{28}}$} $\mathsf{E}_{6}$ $\{1,6\}$ $\mathsf{A}_{2}$ $\mathsf{D}_{4}$ 88 11
$\mathsf{G}_{2,2}^{0}$ $\mathsf{G}_{2}$ $\{1_{\mathrm{s}},2_{\mathrm{l}}\}$ $\mathsf{G}_{2}$ $\varnothing$ 11, 11 11
$^{3}\!{\mathsf{D}_{4,2}^{2}}$}, $^{6}\!{\mathsf{D}_{4,2}^{2}}$} $\mathsf{D}_{4}$ $\{1_{\mathrm{s}},2_{\mathrm{l}},3_{\mathrm{s}},4_{\mathrm{s}}\}$ $\mathsf{G}_{2}$ $\varnothing$ 33, 11 66
$^{1}\!{\mathsf{E}_{6,2}^{16}}$}, $^{2}\!{\mathsf{E}_{6,2}^{16^{\prime\prime}}}$} $\mathsf{E}_{6}$ $\{2_{\mathrm{l}},4_{\mathrm{s}}\}$ $\mathsf{G}_{2}$ $2\mathsf{A}_{2}$ 99, 11 22
$\mathsf{E}_{8,2}^{78}$ $\mathsf{E}_{8}$ $\{7_{\mathrm{s}},8_{\mathrm{l}}\}$ $\mathsf{G}_{2}$ $\mathsf{E}_{6}$ $27$, 11 11
$^{2}\!{\mathsf{E}_{6,2}^{16^{\prime}}}$} $\mathsf{E}_{6}$ $\{1_{\mathrm{us}},2_{\mathrm{s}},6_{\mathrm{us}}\}$ $\mathsf{BC}_{2}$ $\mathsf{A}_{3}$ 88, 66, 11 22
$\mathsf{E}_{7,2}^{31}$ $\mathsf{E}_{7}$ $\{1_{\mathrm{s}},6_{\mathrm{us}}\}$ $\mathsf{BC}_{2}$ $\mathsf{A}_{1}+\mathsf{D}_{4}$ $16$, 88, 11 11
$\mathsf{E}_{8,2}^{66}$ $\mathsf{E}_{8}$ $\{1_{\mathrm{us}},8_{\mathrm{s}}\}$ $\mathsf{BC}_{2}$ $\mathsf{D}_{6}$ $32$, $12$, 11 11
$\mathsf{E}_{7,3}^{28}$ $\mathsf{E}_{7}$ $\{1_{\mathrm{s}},6_{\mathrm{s}},7_{\mathrm{l}}\}$ $\mathsf{C}_{3}$ $\mathsf{D}_{4}$ 88, 11 11
$\mathsf{F}_{4,4}^{0}$ $\mathsf{F}_{4}$ $\{1_{\mathrm{l}},2_{\mathrm{l}},3_{\mathrm{s}},4_{\mathrm{s}}\}$ $\mathsf{F}_{4}$ $\varnothing$ 11, 11 11
$^{2}\!{\mathsf{E}_{6,4}^{2}}$} $\mathsf{E}_{6}$ $\{1_{\mathrm{s}},2_{\mathrm{l}},3_{\mathrm{s}},4_{\mathrm{l}},5_{\mathrm{s}},6_{\mathrm{s}}\}$ $\mathsf{F}_{4}$ $\varnothing$ 22, 11 22
$\mathsf{E}_{7,4}^{9}$ $\mathsf{E}_{7}$ $\{1_{\mathrm{l}},3_{\mathrm{l}},4_{\mathrm{s}},6_{\mathrm{s}}\}$ $\mathsf{F}_{4}$ $3\mathsf{A}_{1}$ 44, 11 11
$\mathsf{E}_{8,4}^{28}$ $\mathsf{E}_{8}$ $\{1_{\mathrm{s}},6_{\mathrm{s}},7_{\mathrm{l}},8_{\mathrm{l}}\}$ $\mathsf{F}_{4}$ $\mathsf{D}_{4}$ 88, 11 11
$^{1}\!{\mathsf{E}_{6,6}^{0}}$} $\mathsf{E}_{6}$ $\{1,2,3,4,5,6\}$ $\mathsf{E}_{6}$ $\varnothing$ 11 11
$\mathsf{E}_{7,7}^{0}$ $\mathsf{E}_{7}$ $\{1,2,3,4,5,6,7\}$ $\mathsf{E}_{7}$ $\varnothing$ 11 11
$\mathsf{E}_{8,8}^{0}$ $\mathsf{E}_{8}$ $\{1,2,3,4,5,6,7,8\}$ $\mathsf{E}_{8}$ $\varnothing$ 11 11
Finally, we give the complete list of isomorphisms and equivalences between isotropic irreducible Tits indices. Two such indices are called isomorphic (denoted by $\cong$) if they differ by an automorphism of the Dynkin diagram of $\widetilde{\Phi}$ and, possibly, by its renaming via an exceptional isomorphism. We say that two indices are equivalent (denoted by $\sim$) if the induced maps $u$ differ by isomorphisms between the corresponding root systems $\widetilde{\Phi}$ and $\Phi$. Equivalent Tits indices are indistinguishable for our purposes.
${{}^{1}\!{\mathsf{A}_{1,1}^{(1)}}}$ $\cong\mathsf{B}_{1,1}\cong\mathsf{C}_{1,1}^{(1)},$ ${{}^{1}\!{\mathsf{D}_{4,1}^{(1)}}}$ $\cong{{}^{1}\!{\mathsf{D}_{4,1}^{(4)}}}\sim{{}^{2}\!{\mathsf{D}_{4,1}^{(1)}}},$
$\mathsf{B}_{2,1}$ $\cong\mathsf{C}_{2,1}^{(2)},$ ${{}^{1}\!{\mathsf{D}_{4,1}^{(2)}}}$ $\sim{{}^{2}\!{\mathsf{D}_{4,1}^{(2)}}}\sim{{}^{3}\!{\mathsf{D}_{4,1}^{9}}}\sim{{}^{6}\!{\mathsf{D}_{4,1}^{9}}},$
$\mathsf{B}_{2,2}$ $\cong\mathsf{C}_{2,2}^{(1)},$ ${{}^{1}\!{\mathsf{D}_{4,2}^{(1)}}}$ $\cong{{}^{1}\!{\mathsf{D}_{4,2}^{(2)}}}\sim{{}^{2}\!{\mathsf{D}_{4,2}^{(1)}}},$
${{}^{2}\!{\mathsf{A}_{3,1}^{(1)}}}$ $\cong{{}^{2}\!{\mathsf{D}_{3,1}^{(2)}}},$ ${{}^{3}\!{\mathsf{D}_{4,2}^{2}}}$ $\sim{{}^{6}\!{\mathsf{D}_{4,2}^{2}}},$
${{}^{2}\!{\mathsf{A}_{3,1}^{(2)}}}$ $\cong{{}^{2}\!{\mathsf{D}_{3,1}^{(1)}}}\sim{{}^{1}\!{\mathsf{A}_{3,1}^{(2)}}}\cong{{}^{1}\!{\mathsf{D}_{3,1}^{(1)}}},$ ${{}^{1}\!{\mathsf{E}_{6,2}^{16}}}$ $\sim{{}^{2}\!{\mathsf{E}_{6,2}^{16^{\prime\prime}}}},$
${{}^{2}\!{\mathsf{A}_{3,2}^{(1)}}}$ $\cong{{}^{2}\!{\mathsf{D}_{3,2}^{(1)}}},$ ${{}^{1}\!{\mathsf{A}_{n,1}^{(d)}}}$ $\sim{{}^{2}\!{\mathsf{A}_{n,1}^{(d)}}}\text{ for }d\geq 3,\,n\geq 5,\,2d=n+1,$
${{}^{1}\!{\mathsf{A}_{3,3}^{(1)}}}$ $\cong{{}^{1}\!{\mathsf{D}_{3,3}^{(1)}}},$ ${{}^{1}\!{\mathsf{D}_{n,r}^{(d)}}}$ $\sim{{}^{2}\!{\mathsf{D}_{n,r}^{(d)}}}\text{ for }d=2^{k}\mid 2n,\,d\geq 1,\,r\geq 1,\,n\geq 5,\,rd\leq n-2.$
@@ 3 Technical lemmas
Note that any parabolic set is saturated (and every saturated set is an intersection of parabolic ones), but $\mathsf{A}_{2}\subseteq\mathsf{G}_{2}$ and $\mathsf{C}_{2}\subseteq\mathsf{BC}_{2}$ are not saturated. Any closed subset $\Sigma$ admits a unique decomposition $\Sigma=\Sigma_{\mathrm{r}}\sqcup\Sigma_{\mathrm{u}}$, where $\Sigma_{\mathrm{r}}=\Sigma\cap(-\Sigma)$ is a closed root subsystem and $\Sigma_{\mathrm{u}}=\Sigma\setminus(-\Sigma)$ is a unipotent set. Conversely, if $\Sigma_{\mathrm{r}}$ is a closed root subsystem, $\Sigma_{\mathrm{u}}$ is a unipotent set disjoint with $\Sigma_{\mathrm{r}}$, and $(\Sigma_{\mathrm{r}}+\Sigma_{\mathrm{u}})\cap\Phi\subseteq\Sigma_{\mathrm{u}}$, then $\Sigma_{\mathrm{r}}\cup\Sigma_{\mathrm{u}}$ is a closed set. A closed set $\Sigma$ is parabolic if and only if there are $w\in\mathrm{W}(\Phi)$ and $J\subseteq\Delta$ such that $w\Sigma=(\Phi^{+}+\mathbb{Z}J)\cap\Phi$. Here $\Phi^{+}\subseteq\Phi$ is the set of positive roots and $\Delta\subseteq\Phi^{+}$ is the set of basic roots. The classes of closed sets, unipotent sets, closed root subsystems, and saturated sets are closed under intersection.
@@
If $\Sigma\subseteq\Phi$ is a closed root subsystem, then the smallest closed set containing $u^{-1}(\Sigma)$ is also a closed root subsystem.
For any closed subset $\Sigma\subseteq\Phi$ let $G_{\Sigma}\leq G$ and $G^{0}_{\Sigma}\leq G$ be the group subsheaves generated by $\bigcup_{\alpha\in\Sigma}U_{\alpha}$ and $L\cup\bigcup_{\alpha\in\Sigma}U_{\alpha}$, they are actually smooth closed group subschemes with connected geometric fibers over $K$ and $G^{0}_{\Sigma}$ is of type (RC) [13, §XXII.5]. More precisely, if $\Sigma$ is a root subsystem, then $G_{\Sigma}$ and $G^{0}_{\Sigma}$ are reductive group subschemes (with induced isotropic structure). If $\Sigma$ is unipotent, then $G_{\Sigma}$ is a unipotent subgroup, i.e. it is an étale twisted form of some $\mathbb{A}^{N}$ as a scheme. Finally, in general $G_{\Sigma}=G_{\Sigma_{\mathrm{r}}}\rtimes G_{\Sigma_{\mathrm{u}}}$ and $G^{0}_{\Sigma}=G^{0}_{\Sigma_{\mathrm{r}}}\rtimes G_{\Sigma_{\mathrm{u}}}$. If $\Sigma$ is parabolic, then $G^{0}_{\Sigma}$ is a parabolic group subscheme. For any parabolic $\Sigma\subseteq\Phi$ the multiplication morphism
@@
$G^{0}_{\Sigma}(K)=G_{\Sigma_{\mathrm{r}}\cap\Phi^{+}}(K)\,G_{\Sigma\cap\Phi^{-}}(K)\,G_{\Sigma\cap\Phi^{+}}(K)\,L(K).$
@@
We may assume that $\operatorname{Spec}(K)$ is connected. In this case there is a maximal isotropic structure $(L^{\prime},G^{\prime}_{\alpha})_{\alpha\in\Phi^{\prime}}$ and a map $v\colon\Phi^{\prime}\to\Phi\cup\{0\}$ such that $\Phi\subseteq v(\Phi^{\prime})$, $G^{\prime}_{\alpha}=G_{v^{-1}(\alpha)}$, $L^{\prime}=G_{\Phi\cap v^{-1}(0)}$ [13, XXVI.7.4.2]. Gauss decomposition holds for $G$ with respect to a maximal isotropic structure by [13, corollary XXVI.5.2], i.e.
$G(K)=G^{\prime}_{\Phi^{\prime+}}(K)\,G^{\prime}_{\Phi^{\prime-}}(K)\,G^{\prime}_{\Phi^{\prime+}}(K)\,L^{\prime}(K).$
@@
$G^{0}_{\Sigma}(K)=G_{\Sigma\cap\Phi^{+}}(K)\,G_{\Sigma\cap\Phi^{-}}(K)\,G_{\Sigma\cap\Phi^{+}}(K)\,L(K)$
@@
$G^{0}_{\Sigma}(K)$ $=G_{\Sigma_{\mathrm{r}}\cap\Phi^{+}}(K)\,G_{\Sigma_{\mathrm{r}}\cap\Phi^{-}}(K)\,G_{\Sigma_{\mathrm{u}}}(K)\,G_{\Sigma_{\mathrm{r}}\cap\Phi^{+}}(K)\,L(K)$
$=G_{\Sigma_{\mathrm{r}}\cap\Phi^{+}}(K)\,G_{\Sigma\cap\Phi^{-}}(K)\,G_{\Sigma\cap\Phi^{+}}(K)\,L(K).\qed$
For any closed subsets $\Sigma,\Sigma^{\prime}\subseteq\Phi$ the intersection $G^{0}_{\Sigma}\cap G^{0}_{\Sigma^{\prime}}$ is a smooth closed subscheme and its fiberwise connected component is $G^{0}_{\Sigma\cap\Sigma^{\prime}}$ by [13, proposition XXII.5.4.5].
@@
For the first claim it suffices to prove that $G^{0}_{\Sigma}\cap G^{0}_{\Sigma^{\prime}}=G^{0}_{\Sigma\cap\Sigma^{\prime}}$ if $\Sigma^{\prime}\supseteq\Phi^{+}$ is parabolic and $K$ is local. Take $g\in G^{0}_{\Sigma}(K)\cap G^{0}_{\Sigma^{\prime}}(K)$. By lemma 1 we can multiply $g$ from both sides by elements of $G^{0}_{\Sigma\cap\Sigma^{\prime}}$ to get a new element $g^{\prime}\in G_{\Sigma\cap\Phi^{-}}(K)\cap G^{0}_{\Sigma^{\prime}}(K)$. But such $g^{\prime}$ necessarily lies in $G_{\Sigma\cap\Sigma^{\prime}\cap\Phi^{-}}$.
Now let us prove the second claim. Without loss of generality $\Sigma^{\prime}\subseteq\Phi^{+}$, so
$G^{0}_{\Sigma}\cap G_{\Sigma^{\prime}}=G^{0}_{\Sigma}\cap G^{0}_{\Phi^{+}}\cap G_{\Sigma^{\prime}}=G^{0}_{\Sigma\cap\Phi^{+}}\cap G_{\Sigma^{\prime}}=G_{\Sigma\cap\Phi^{+}\cap\Sigma^{\prime}}=G_{\Sigma\cap\Sigma^{\prime}}.\qed$
Recall [25, §4] that a $2$-step $K$-module $(M,M_{0})$ consists of
@@
  • a central subgroup $M_{0}\leq M$ such that $M,M]^{\cdot}\leq M_{0}$;
@@
  • a right action $({-})\cdot({=})\colon M\times K\to M$ of the multiplicative monoid by group automorphisms such that $m\cdot k,m^{\prime}\cdot k^{\prime}]^{\cdot}=kk^{\prime}[m,m^{\prime}]^{\cdot}$, $m\cdot(k+k^{\prime})=m\cdot k\dotplus kk^{\prime}\tau(m)\dotplus m\cdot k^{\prime}$ for some (uniquely determined) $\tau(m)\in M_{0}$, $m_{0}\cdot k=k^{2}m_{0}$ for $m_{0}\in M_{0}$.
@@
$\tau(\mathbin{\dot{-}}m)$ $=\mathbin{\dot{-}}\tau(m),$ $\tau(m\cdot k)$ $=k^{2}\tau(m),$